OpenStudy (anonymous):

quadratic formula for x^2-7x+2=0

jimthompson5910 (jim_thompson5910):

\[\Large x=\frac{-B\pm\sqrt{B^2-4AC}}{2A}\] \[\Large x=\frac{-\left(-7\right)\pm\sqrt{\left(-7\right)^2-4\left(1\right)\left(2\right)}}{2\left(1\right)}\] \[\Large x=\frac{7\pm\sqrt{49-8}}{2}\] \[\Large x=\frac{7\pm\sqrt{41}}{2}\] \[\Large x=\frac{7\pm\sqrt{41}}{2}\] \[\Large x=\frac{7+\sqrt{41}}{2} \ \text{or} \ x=\frac{7-\sqrt{41}}{2}\]

jimthompson5910 (jim_thompson5910):

first line you have should be +49/4 on the end

jimthompson5910 (jim_thompson5910):

yeah you caught it

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np, probably why the quad formula is nicer to work with